40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 26 17:45:21 2014, Steve, Update, VAC, vac pressure rose to 1.3 mTorr noN2.pngbacktoVacNormal.png
    Reply  Sun Dec 28 07:29:16 2014, Steve, Update, VAC, vacuum is normal 3d.png
    Reply  Tue Feb 3 15:20:13 2015, Steve, Koji, Update, VAC, Pneumatic pressure is being read again N2pressureReadingIsBack.png
    Reply  Sun Mar 22 18:42:40 2015, Steve, Update, VAC, vac pressure rose to 1.2mTorr NoN2.pngPressureReadingWorks.png
       Reply  Mon Mar 23 10:36:55 2015, ericq, Update, VAC, Pressure watch script 
          Reply  Fri Apr 24 17:30:32 2015, Jenne, Update, VAC, Pressure watch script broken 
             Reply  Sat Apr 25 18:50:47 2015, ericq, Update, VAC, Pressure watch script broken 
                Reply  Sat Apr 25 22:17:49 2015, rana, Update, CDS, MXstream restart script working (beta) 
                   Reply  Tue May 19 17:03:34 2015, Koji, Update, CDS, MXstream restart script working (beta) 
                      Reply  Tue May 19 18:38:33 2015, rana, Update, CDS, MXstream restart script working (beta) 
          Reply  Mon Apr 27 14:37:55 2015, Steve, Update, VAC, N2 pneumatic pressure watch  
             Reply  Thu Apr 30 16:30:25 2015, Steve, Update, VAC, N2 pneumatic pressure watch set up 
          Reply  Tue Jan 31 14:14:29 2017, rana, Update, VAC, Pressure watch script 
       Reply  Sun May 15 19:48:08 2016, Steve, Update, VAC, run out of N2 noN2.png
          Reply  Tue Jun 14 10:13:34 2016, Steve, Update, VAC, temporary N2 supply line afterN2work.png
             Reply  Fri Jun 17 15:22:31 2016, gautam, Update, VAC, N2 supply line restored after retiling c0vac_06172016.png
Message ID: 10968     Entry time: Tue Feb 3 15:20:13 2015     In reply to: 10843
Author: Steve, Koji 
Type: Update 
Category: VAC 
Subject: Pneumatic pressure is being read again 

[Koji, Steve]

Summary

The N2 pressure reading (C1:VAC-N2PRES) is now up-to-date after rebooting c1vac1.
The vaccum system is "Vacuum normal". We now have a space pressure transducer.

Introduction

Our vacuum valves are manipulated with 60~75 PSI of nitrogen. All the valves are configured to be closed in the case of low N2 supply pressure.
In order to avoid this safety shutdown accidentally triggered, we have two N2 cylinders to sustain the vacuum valves. When one cylinder goes to low
the mechanical valve switches over to the other cylinder.

We have the monitor channel for this (combined) cylinder pressure. One shoulbe be able to see periodical pressure variation when the auto cylinder
switch is operating. However, the nirogen pressure reading got stuck at 66 PSI on Dec.16, 2014 (See attached 60-day plot of N2 supply pressure). 

What we did

This morning we tracked down the cause of the trouble. We first closed the valves on EPICS and started to vary the N2 pressure.

Our first guess was the pressure transducer (Omega #236PC100GW) that was already 15 yrs old. We even has a sensor spare for replacement.
But it turned out that the direct voltage reading (to be 1mV/PSI) is changing correctly. The second guess was Omega Controller-Monitor
#DPiS32-C24 that is reading the voltage from the tranceducer. The display on this small black unit was changing corresponding to the
pressure change.

So our thought was
1) RS232C of the monitor unit is not working correctly
or
2) c1vac1 is not communicating with the monitor unit.

We wondered what could cause c1vac1 not communicating with the monitor unit, but we were afraid that some function got stuck
during either the nodus upgrade or chiara rebooting (or something else). So we decided to reboot c1vac1

In order to avoid any glitch in the main vacuum pressure, Steve disconnected some of the controller connectors for the closed valves.
We did this treatment before and it was successful.

Then c1vac1 was rebooted just by telnet and type reboot in the terminal.
Once the target is back in action, we noticed that the monitor value started to move.

Steve reverted the cables to the valves and operated the valves to recover "Vacuum Normal" state. Everything is now nicely settled.

Attachment 1: N2pressureReadingIsBack.png  119 kB  | Hide | Hide all
N2pressureReadingIsBack.png
ELOG V3.1.3-